OpenStudy (anonymous):

9^n=49 for solving exponential equations with logarithms

OpenStudy (bemoseemo):

Let's solve your equation step-by-step. 9n=49 Solve Exponent. 9n=49 log(9n)=log(49)(Take log of both sides) n*(log(9))=log(49) n= log(49) log(9) n=1.771244 Answer: n=1.771244

OpenStudy (bemoseemo):

@kendrick1738